# Diffusiophoretic corner flows

### v1

@@ IV Phoretic walls with uniform coverage
To study the effect of chemical activity on corner flows, we first focus on the case when the coverage by catalyst is uniform, and the heterogeneity of the concentration field stems purely from the geometry of the wedge. In the two examples studied in the following, either one or both walls are active, and the problem admits analytical solutions.
@@ IV.1 A single active wall
$c(\rho,\theta)=-\frac{A\rho\sin(\theta-\alpha)}{\cos\alpha},$ (23)
from which the tangential gradient dictates the slip velocity distribution on the surface as
@@
with the boundary conditions $f(0)=0$, $f^{\prime}(0)=1$, $f(\alpha)=0$, $f^{\prime}(\alpha)=0$. This is a particular form of the general solution to planar elasticity problems for the biharmonic Airy stress function due to [38]. The resulting stream function reads:
@@
The resulting flow fields are drawn in figure 2 for three values of the wedge opening angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The flow is asymmetric, reflecting the boundary conditions.
Figure 2: Phoretic flow induced by the activity of one phoretic wall in a wedge of angle $\alpha$ for (left to right) $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. Flow streamlines are marked in white. The colour map indicates the total velocity magnitude.
@@ IV.2 Two phoretic walls
Equation (27) remains valid with the new boundary conditions $f(0)=f(\alpha)=0$ and $f^{\prime}(0)=f^{\prime}(\alpha)=1$. The stream function in this case takes on the form
@@
Figure 3: Phoretic flow induced by the activity of two phoretic walls in a wedge of angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$.
The resulting flow fields are drawn in figure 3 for three values of the wedge opening angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The flow is symmetrical, with the fluid being dragged in from the sides and pushed away along the bisector direction.
@@ V The diffusion problem with an active and absorbing wall
To explore the flow in a more general case, we focus on a corner with a single active sector of length $\ell=b-a$ placed on one of the boundaries at a distance $a$ (between $\rho=a$ and $\rho=b$), as depicted in figure 1. The problem then, contrary to the previous cases, has two natural length scales: $\ell$ and $a$.
The concentration field satisfies the Laplace’s equation, $\nabla^{2}c=0$, with the boundary conditions
$\frac{\partial c}{\partial\theta}=-A\rho\qquad$ $\text{for $\theta=0$ and $a<\rho<b$},$ (32)
$c=0\qquad$ $\text{on other boundaries}.$ (33)
V.1 Solution of the diffusion problem for a single catalytic segment
We consider a single catalytic segment $[a,b]$ placed on one of the walls ($\theta=0$), as sketched in figure 1, and analyse the solution in several specific cases.
V.1.1 The case of $a\neq 0$ and $b<\infty$
Let the wall with $\theta=\alpha$ be ideally absorptive. Then, the boundary conditions take the form
$\frac{\partial c}{\partial\theta}$ $=-A\rho\mathds{1}_{[a,b]}$ $\textrm{for}\quad\theta=0$ (34a)
$c$ $=0$ $\textrm{for}\quad\theta=\alpha.$ (34b)
It is assumed that $0<\alpha<2\pi$. The solution in Mellin space is
@@ V.1.1 The case of $a\neq 0$ and $b<\infty$
It has singularities at $p^{\prime}_{k}$ such that $p^{\prime}_{k}\alpha=\frac{\pi}{2}+k\pi$ for $k\in\mathbb{Z}$, and also at $p_{0}=-1$, and a removable singularity at $p=0$.
The concentration is then written as $c=I_{b}-I_{a}$, where
$I_{a}=\frac{1}{2\pi i}\int_{-i\infty}^{i\infty}A\frac{a}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}\left(\frac{a}{\rho}\right)^{p}dp.$ (36)
For $a\leq\rho$, the integral can be evaluated by taking a square contour over the half-plane $\text{Re}(p)>0$; for $a>\rho$, the contour is taken over $\text{Re}(p)<0$. The choice ensures that the integral over three sides of the square contour vanishes.
The poles are listed as
@@
Note that the residue at p=1p=-1 is responsible for satisfying the condition (34a). In the special cases of $\alpha=\frac{\pi}{2},\pi,\frac{3\pi}{2}$ solutions can be obtained by taking, for example, the limit of $\alpha\to\frac{\pi}{2}$. The solutions for $\alpha=\frac{\pi}{2}-\epsilon$ and $\alpha=\frac{\pi}{2}+\epsilon$ can then be compared to check if the desired accuracy has been achieved for a chosen $\epsilon$.
V.1.2 The case of $a=0$, $b<\infty$ i.e. sector $[0,b]$
@@
V.1.3 The case of $a\neq 0$ and $b=\infty$ i.e. sector $a,\infty[$
In this case existence of Mellin transform requires that $\text{Re}(p)\leq-1$.
@@
Figure 4: Solute concentration fields for an ideally absorptive wall at $\theta=\pi/6$, a phoretic wall at $\theta=0$ with a catalytic sector at $(a,b)=\{(1,3),(0,3),(3,\infty)\}$ marked in red. Here, we assumed $A=1$ and the plotted radius of the wedge is $\rho<4$. The scale bar for the absolute concentration field, $|c|$, is common for plots (a), (b) and different for panel (c).
@@ VI.1 Solution for a single catalytic sector with reflective walls
If $c_{\alpha}$ is a solution for the active sector with activity $A$ on the wall $\theta=\alpha$ i.e., for boundary conditions
@@
The choice of $\gamma$ such that $\textrm{min}\{p^{\prime}_{k}:p^{\prime}_{k}>0\}=\frac{\pi}{\alpha}>\gamma>0$ guarantees that the Mellin transform of $c_{0}$ exists. Integrating over a square contour in one of the half-planes, we obtain
@@ VI.2 Diffusion for multiple catalytic sectors
An example of a setting involving multiple sectors is presented in figure 5 for two active patches of opposite activity on each of the walls, and for different wedge opening angles. The activity of patches closer to the tip of the wedge is diminished by their mutual influence, and the concentration fields produced by patches further away from the confinement are more pronounced. This demonstrates the intricate interaction between the geometry of the wedge and the chemical activity of its walls.
Figure 5: Solute concentration fields for wedges with multiple active sectors. (a)-(d) The wedge angles are $\alpha=\{4\pi/7,3\pi/7,5\pi/6,\pi/6\}$. There are two active sectors of opposite activity $|A|=1$ on each wall. They span the sections $\rho\in(0.5,1.5)$ and $\rho\in(2.5,3.5)$. The plotted radius of the wedge is $\rho<4$. Patches of positive activity are marked in red, while those of negative activity are marked in blue. Panels (a)-(c) share the scale bar placed in the middle of the figure, while the scale bar for panel (d) is in the top right corner.
@@ VI.3 The diffusion problem for an arbitrary catalytic wall
Let us now consider a setting where the wall at $\theta=0$ has arbitrary activity $A=A(\rho)$ and the wall at $\theta=\alpha$ is reflective. We will approach this problem using the Green’s function, because the Stokes equation is linear. Let $c^{G}_{0}(\rho,\rho^{\prime},\theta)$ be the Green’s function for activity $\delta(\rho-\rho^{\prime})$. Boundary conditions take the following form,
@@
$\bar{c}^{G}_{0}(p,\rho^{\prime},\theta)=-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\rho^{\prime p},$ (54)
and the general solution for the transformed concentration becomes
@@ VII Solution of the biharmonic equation for given solute concentration field
Having resolved the concentration field in various cases and for different boundary conditions on each wall, we now turn to the calculation of the corresponding flow field. The boundary conditions for the stream function $\Psi(\rho,\theta)$, that satisfies the biharmonic equation (12), are
@@
Figure 6: (a) Isolines of the stream function $\psi$ of diffusiophoretic corner flow for $\theta=\pi/3$, with active sectors at $\rho\in(0,1)$, emitting at $\theta=0$ and absorbing at $\theta=\alpha$. The flow in the corner eddy (yellow) is counter-clockwise and drives clockwise rotation of another eddy further away from the corner. Colours code the magnitude of $\psi$. (b) Transversal velocity profile $v_{\theta}(\rho)$, on the bisector angle of the wedge. Roots of the velocity indicate the centres of vortices. The slight wiggle for small $\rho$ is likely a numerical artifact.
VIII The diffusion problem with analytic Neumann boundary conditions
Here, we present an auxiliary result for the diffusion problem when the activity patterns are given by analytic functions, rather than step-like profiles discussed earlier. Let $A^{1}(\rho)$ and $A^{2}(\rho)$ represent the (analytic) distribution of activity on the walls with $\theta=0$ and $\theta=\alpha$, respectively. We make the following ansatz for the concentration profile $c$ in the fluid domain
@@ VIII The diffusion problem with analytic Neumann boundary conditions
and expand $A^{1}$ and $A^{2}$ in power series in $\rho$ as
$A^{1}=\sum_{n\geq 1}A^{1}_{n}\rho^{n},\ A^{2}=\sum_{n\geq 1}A^{2}_{n}\rho^{n}.$ (65a)
@@
$F_{n}$ $=-\frac{A^{1}_{n-1}\cos(n\alpha)+A^{2}_{n-1}}{n\sin(n\alpha)},$ (66a)
$G_{n}$ $=-\frac{A^{1}_{n-1}}{n},$ (66b)
@@
IX Conclusions
In this paper, inspired by Moffatt eddies [25, 26] that emerge in wedge-like geometry under boundary forcing, we have shown that corner geometry with phoretically active sectors can be used to generate similar eddies. Our solution involved the exact solution of the diffusion problem in the wedge geometry using the Mellin transform formalism, followed by a solution of the biharmonic equation for the stream function of the associated flow. The proposed diffusiophoretic framework relies on a fixed-flux boundary condition for the solute release from a catalytic patch, which is appropriate when the diffusion of the solute is much faster than its advection (that is, in the limit of low Damköhler number).
In simple cases of uniform coverage of the walls with chemical activity, the geometric asymmetry alone is responsible for heterogeneous distribution of solute concentration, which drives the flow. In this case, the problem admits a simple analytical solution akin to the classical Taylor’s scraper problem when one wall is chemically active. Similarly, an analytical solution is also available when both walls are covered with catalyst. These solutions can serve as benchmarks for numerical computations, particularly when solving for the concentration field with non-trivial phoretic boundary conditions.
@@ IX Conclusions
Data availability statement The Mathematica notebooks used to compute the results of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.16938259, reference number 16938259.

### v2

@@ IV Phoretic walls with uniform coverage
To study the effect of chemical activity on corner flows, we first focus on the case when the coverage by catalyst is uniform, and the heterogeneity of the concentration field stems purely from the geometry of the wedge. In the two examples studied in the following, either one or both walls are active, and the problem admits analytical solutions. In this case, the problem has no natural length scale, and the velocity scale is given purely by the activity and mobility at the surface.
@@ IV.1 A single active wall
$c(\rho,\theta)=-\frac{A\rho\sin(\theta-\alpha)}{\cos\alpha}.$ (24)
The fact that the concentration increases linearly with the distance from the tip is a consequence of the fixed flux boundary condition and the fact that the active patch extends to infinity. In reality, the finite size of the active patch limits the concentration field, as we discuss later on. As a result, the tangential gradient of the concentration profile yields a constant slip velocity distribution on the surface,
@@
with the boundary conditions $f(0)=0$, $f^{\prime}(0)=1$, $f(\alpha)=0$, $f^{\prime}(\alpha)=0$. We note that by Eq. (14) the boundary conditions for $f$ correspond to the normal velocity at the surface, while those for $f^{\prime}$ pertain to the induced slip (tangential) velocity component. This is a particular form of the general solution to planar elasticity problems for the biharmonic Airy stress function due to [72]. The resulting stream function reads:
@@
The resulting flow fields are drawn in figure 2 for three values of the wedge opening angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The flow is asymmetric, reflecting the boundary conditions. We observe the strongest flow along the active boundary, and a gradual decay of the velocity field closer to the other, no-slip wall. We note that the magnitude of the bulk velocity field remains a large fraction of the surface slip velocity, confirming the pronounced effect of surface actuation.
Figure 2: Diffusioosmotic flow induced by the activity of one phoretic wall in a wedge of angle $\alpha$ for (left to right) $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. Flow streamlines are marked in white. The colour map indicates the total velocity magnitude. The emergent bulk flow remains comparable in magnitude to the driving slip flow on tha active boundary, and decays rapidly close to the inert, no-slip wall.
@@ IV.2 Two phoretic walls
Equation (28) remains valid with the new boundary conditions $f(0)=f(\alpha)=0$ (no normal velocity at the surface) and $f^{\prime}(0)=f^{\prime}(\alpha)=1$ (two active walls with non-zero slip velocity). The stream function in this case takes on the form
@@
Figure 3: Diffusioosmotic flow induced by the activity of two phoretic walls in a wedge of angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The colour map indicates the total velocity magnitude. We note a strong flow close to the driving active boundaries, and a counterflow along the wedge bisector.
The resulting flow fields are drawn in figure 3 for three values of the wedge opening angle $\alpha=\{\pi/4,\pi/2,3\pi/4\}$. The flow is symmetrical, with the fluid dragged along the walls towards the wedge vertex and expelled along the bisector.
@@ V The diffusion problem with an active and absorbing wall
To explore the flow in a more general case, we focus on a corner with a single active sector of length $\ell=b-a$ placed on one of the boundaries at a distance $a$ (between $\rho=a$ and $\rho=b$), as depicted in figure 1. The other wall is perfectly absorbing. Contrary to the previous cases, the problem now has two natural length scales: $\ell$ and $a$.
The concentration field satisfies the Laplace equation, $\nabla^{2}c=0$, with the boundary conditions
$\frac{\partial c}{\partial\theta}=-A\rho\mathds{1}_{[a,b]}\quad$ $\textrm{for}\quad\theta=0,$ (33a)
$c=0\hskip 42.67912pt$ $\textrm{for}\quad\theta=\alpha,$ (33b)
where $\mathds{1}_{[a,b]}$ is the indicator function of set $[a,b]$.
V.1 The case of $a\neq 0$ and $b<\infty$
The solution in Mellin space is
@@ V.1 The case of $a\neq 0$ and $b<\infty$
It is assumed that $0<\alpha<2\pi$. The solution has singularities at $p^{\prime}_{k}$ such that $p^{\prime}_{k}\alpha=\frac{\pi}{2}+k\pi$ for $k\in\mathbb{Z}$, and also at $p_{0}=-1$, and a removable singularity at $p=0$.
To evaluate the concentration field, we must now invert the Mellin transform. The concentration in real space can be written as $c=I_{b}-I_{a}$, where
$I_{a}=\frac{1}{2\pi i}\int_{\gamma-i\infty}^{\gamma+i\infty}A\frac{a}{p(p+1)}\frac{\sin(p(\alpha-\theta))}{\cos p\alpha}\left(\frac{a}{\rho}\right)^{p}dp.$ (35)
The integral can be evaluated by applying Cauchy’s residue theorem to a square integration contour sketched in figure 4. For $a\leq\rho$, the integration contour lies in the half-plane $\text{Real}(p)>\gamma$. For $a>\rho$, the appropriate contour lies in the $\text{Real}(p)<\gamma$ half-plane. Such a choice ensures that the integral over three sides of the square contour vanishes when the contour is extended to infinity.
Figure 4: Contours of integration for the evaluation of the inverse Mellin transform. The green contour is used for $a>\rho$, and the blue contour for $a\leq\rho$. Both integration contours are shifted by $\gamma$ along the real axis, and the poles of the integrand are denoted by $p_{k}$, where $k\in\mathbb{Z}$. To evaluate the integral, one takes the limit of the square side length approaching infinity. In the limit, contributions from the three dashed sides of each square contour vanish, and the desired integral along the imaginary axis can be evaluated using the method of residues.
In particular, for the integral in Eq. (35), we select $\gamma=0$. The poles are listed as
@@
Note that the residue at p=1p=-1 is responsible for satisfying the condition (33a). In the special cases of $\alpha={\pi}/{2},\pi,{3\pi}/{2}$ solutions can be obtained by taking, for example, the limit of $\alpha\to{\pi}/{2}$. The solutions for $\alpha={\pi}/{2}-\epsilon$ and $\alpha={\pi}/{2}+\epsilon$ can then be compared to check if the desired accuracy has been achieved for a chosen $\epsilon$.
V.2 The case of $a=0$, $b<\infty$ i.e. sector $[0,b]$
@@
V.3 The case of $a\neq 0$ and $b=\infty$ i.e. sector $a,\infty[$
In this case existence of Mellin transform requires that $\gamma\leq-1$.
@@
Figure 5: Solute concentration fields, (a)–(c), and isolines of the stream function $\psi$, (d)–(f), for an ideally absorptive wall at $\theta=\pi/6$, a phoretic wall at $\theta=0$ with a catalytic sector at $(a,b)=\{(1,3),(0,3),(1,\infty)\}$ marked in red. Here, we assumed $A=1$ and the plotted radius of the wedge is $\rho<4$. The scale bar for the absolute concentration field, $|c|$, is common for plots (a), (b) and different for panel (c).
@@ VI.1 Solution for a single catalytic sector with reflective walls
If $c_{\alpha}$ is a solution for the active sector with activity $A$ on the wall $\theta=\alpha$ i.e., for the boundary conditions
@@
The choice of $\gamma$ such that $\textrm{min}\{p^{\prime}_{k}:p^{\prime}_{k}>0\}={\pi}/{\alpha}>\gamma>0$ guarantees that the Mellin transform of $c_{0}$ exists. Integrating over a square contour in one of the half-planes, we obtain
@@ VI.2 Diffusion for multiple catalytic sectors
An example of a setting involving multiple sectors is presented in figure 6 for two active patches of opposite activity on each of the walls, and for different wedge opening angles. The effect of activity of patches closer to the tip of the wedge is diminished by their mutual influence, and the concentration fields produced by patches further away from the confinement are more pronounced. This demonstrates the intricate interaction between the geometry of the wedge and the chemical activity of its walls.
Figure 6: Solute concentration fields, (a)-(b), and corresponding stream functions $\psi$, (c)-(d), for wedges with multiple active sectors. The wedge angles are $\alpha=\{4\pi/7,3\pi/7\}$. There are two active sectors of opposite activity $|A|=1$ on each wall. They span the sections $\rho\in(0.5,1.5)$ and $\rho\in(2.5,3.5)$. The plotted radius of the wedge is $\rho<4$. Patches of positive activity are marked in red, while those of negative activity are marked in blue. Panels (a)-(b) share the concentration scale bar placed in the middle of the figure, while the stream function scale bar for panels (c)-(d) is in the top right corner.
@@ VI.3 The diffusion problem for an arbitrary catalytic wall
Let us now consider a setting where the wall at $\theta=0$ has arbitrary activity $A=A(\rho)$ and the wall at $\theta=\alpha$ is reflective. We will approach this problem using the Green’s function method [73], because the Stokes equation is linear. Let $c^{G}_{0}(\rho,\rho^{\prime},\theta)$ be the Green’s function, which is a formal solution to the Laplace equation with a point activity $\delta(\rho-\rho^{\prime})$ at $\rho=\rho^{\prime}$, $\theta=0$. Boundary conditions for $c^{G}_{0}(\rho,\rho^{\prime},\theta)$ take the following form,
@@
$\bar{c}^{G}_{0}(p,\rho^{\prime},\theta)=-\frac{\cos(p(\alpha-\theta))}{p\sin(p\alpha)}\rho^{\prime p}.$ (53)
$\bar{c}_{0}(p,\theta)$ is obtained by accounting for all point activities $A(\rho^{\prime})\delta(\rho-\rho^{\prime})$ along the boundary
@@ VII Solution of the biharmonic equation for given solute concentration field
Having resolved the concentration field in various cases and for different boundary conditions on each wall, we now turn to the calculation of the corresponding flow field. The boundary conditions for the stream function $\Psi(\rho,\theta)$, which satisfies the biharmonic equation (13), are
@@
It is useful to consider an experimentally practical scenario of the wedge opening angle of $\alpha=\pi/2$, for which we present the flow field in figure 8. In this case, a single vortical structure emerges, fueled by catalytic reactions on the walls close to the corner. The active patches cover the area of $\rho<1$ on both walls. The associated transversal velocity profile confirms that the vortex does not extend far beyond the active region.
Figure 7: (a) Isolines of the stream function $\psi$ of diffusioosmotic corner flow for $\theta=\pi/3$, with active sectors at $\rho\in(0,1)$, emitting at $\theta=0$ and absorbing at $\theta=\alpha$. The flow in the corner eddy (yellow) is counter-clockwise and drives clockwise rotation of another eddy further away from the corner. Colours code the magnitude of $\psi$. (b) Transversal velocity profile $v_{\theta}(\rho)$, on the bisector angle of the wedge. Roots of the velocity indicate the centres of vortices.
Figure 8: Diffusioosmotic corner flow for the wedge opening angle of $\pi/2$. (a) Isolines of the stream function show a single vortical structure in the corner. The active patches covers both walls close to the tip (for $\rho<1$), with emission at the horizontal wall and absorption at the vertical surface. (b) Transversal velocity profile along the bisector line of $\theta=\pi/4$.
VIII Diffusion with analytic Neumann boundary conditions
Here, we present an auxiliary result for the diffusion problem when the activity patterns are given by analytic functions, rather than step-like profiles discussed earlier. Let $A^{(1)}(\rho)$ and $A^{(2)}(\rho)$ represent the (analytic) distribution of activity on the walls with $\theta=0$ and $\theta=\alpha$, respectively. We make the following ansatz for the concentration profile $c$ in the fluid domain
@@ VIII Diffusion with analytic Neumann boundary conditions
and expand $A^{(1)}$ and $A^{(2)}$ in power series in $\rho$ as
$A^{(1)}=\sum_{n\geq 1}A^{(1)}_{n}\rho^{n},\ A^{(2)}=\sum_{n\geq 1}A^{(2)}_{n}\rho^{n}.$ (64a)
@@
$F_{n}$ $=-\frac{A^{(1)}_{n-1}\cos(n\alpha)+A^{(2)}_{n-1}}{n\sin(n\alpha)},$ (65a)
$G_{n}$ $=-\frac{A^{(1)}_{n-1}}{n},$ (65b)
@@
IX Discussion and conclusions
In this paper, inspired by Moffatt eddies [49, 50] that emerge in wedge-like geometries under mechanical boundary forcing, we have shown that corner geometry with phoretically active sectors can be used to generate similar eddies. Our solution involved the exact solution of the diffusion problem in the wedge geometry using the Mellin transform formalism, followed by a solution of the biharmonic equation for the stream function of the associated flow. The proposed diffusioosmotic framework relies on a fixed-flux boundary condition for the solute release from a catalytic patch, which is appropriate when the diffusion of the solute is much faster than its adsorption (that is, in the limit of low Damköhler number) and that the advection is slow compared to diffusion (corresponding to the limit of low Péclet number).
In the simple case of uniform coverage of the walls with chemical activity, the geometric asymmetry alone is responsible for heterogeneous distribution of solute concentration, which drives the flow. In such a setting, the problem admits a simple analytical solution akin to the classical Taylor’s scraper problem when one wall is chemically active. Similarly, an analytical solution is also available when both walls are covered with catalyst. These solutions can serve as benchmarks for numerical computations, particularly when solving for the concentration field with non-trivial phoretic boundary conditions.
@@ IX Discussion and conclusions
Declaration of interests The authors report no conflict of interest.
Data availability statement The Mathematica notebooks used to compute the results of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.16938258, reference number 16938258.